OpenStudy (anonymous):

could i find the two missing side lengths of a right triangle if i only know one side length and angle measure (other than 90 degrees)?

OpenStudy (anonymous):

pythagorean theorem (it depends on what you're missing..pythagorean theorem is c^2 = a^2 + b^2 where a and b are the legs and c is the hypotenuse) 2. yes you can...you use sin/cos/tan 3. yes you can...you use sin/cos/tan 4. special right triangles are the 30-60-90 and the 45-45 90 in 30-60-90 triangles the legs are always in the form of this: hypotenuse = twice the length of the side opposite 30 side opposite 30 = 1/2 of the hypotenuse side opposite 60 = the length of the side opposite 60 x sqrt 3 in 45-45-90 the sides opposite 45 are equal and they are equal to sqrt 2 times the hypotenuse not sure about the last sentence...but i know that they are equal and there is a sqrt of 2 somewhere...just dunno if it's in the hypotenuse or the legs

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

welcome can u put best response

OpenStudy (anonymous):

plz i

OpenStudy (anonymous):

sure thing